# Pythia 8 and Air Shower Simulations: A Tuning Perspective Chloé Gaudu1

### v1

@@ 2 Pythia 8
A previous implementation of Pythia 8 in Corsika 8 was done for version 8.3.07 using the PythiaCascade class for hadron-ion collision systems [26, 27]. This work follows in the footsteps of previous study, but now focusing on the Angantyr model in Pythia 8.3.12 and its opportunities.
The current implementation of Pythia in its 8.3.12 version, alongside with the Angantyr model, in Corsika 8 is at a preliminary stage. After initializing the shower input setting, Corsika calls either Pythia or Angantyr depending on the collision system at hand (pp or hA/AA). The latest update brings energy and beam switching features into play, allowing a single Pythia instance to run and perform any collisions for a given projectile, target and center-of-mass energy on an event-by-event basis. The inelastic cross-section is interpolated from pre-generated tables, and secondary final state particles are obtained for each collision system. Tracking occurs internally within Corsika, storing all shower events in a particle stack, resulting in Corsika output files offering insights into shower profiles and ground-level particles among other information.
All tests run with version 8.3.12 were targeted at validating proton-proton collisions between Pythia wherin Corsika 8 and standalone Pythia simulations, as well as checking proton primary air showers generation. Before running air shower simulations with nuclear primaries, several issues must be resolved: the long initialization time required for Angantyr, some gaps in the tabulated cross-sections for nuclear projectiles and targets, and the handling of the nuclear remnants. Optimizing the usage of reuse files within Pythia may cut down initialization time of Angantyr, while adopting a semi-superposition model (similarly done for Sibyll 2.3d) could bridge the gaps in the cross-section tables.
@@ 3 Tuning of Pythia 8
[36] | & charged particle production | 250 | p, K+p, pp | ||

### v2

@@ 2 Pythia 8
A previous implementation of Pythia 8 in Corsika 8 was done for version 8.3.07 using the PythiaCascade class for hadron-ion collision systems [26, 27]. This work follows in the footsteps of previous study, but now focusing on the Angantyr model in Pythia 8.312 and its opportunities.
The current implementation of Pythia in its 8.312 version, alongside with the Angantyr model, in Corsika 8 is at a preliminary stage. After initializing the shower input setting, Corsika calls either Pythia or Angantyr depending on the collision system at hand (pp or hA/AA). The latest update brings energy and beam switching features into play, allowing a single Pythia instance to run and perform any collisions for a given projectile, target and center-of-mass energy on an event-by-event basis. The inelastic cross-section is interpolated from pre-generated tables, and secondary final state particles are obtained for each collision system. Tracking occurs internally within Corsika, storing all shower events in a particle stack, resulting in Corsika output files offering insights into shower profiles and ground-level particles among other information.
All tests run with version 8.312 were targeted at validating proton-proton collisions between Pythia wherin Corsika 8 and standalone Pythia simulations, as well as checking proton primary air showers generation. Before running air shower simulations with nuclear primaries, several issues must be resolved: the long initialization time required for Angantyr, some gaps in the tabulated cross-sections for nuclear projectiles and targets, and the handling of the nuclear remnants. Optimizing the usage of reuse files within Pythia may cut down initialization time of Angantyr, while adopting a semi-superposition model (similarly done for Sibyll 2.3d) could bridge the gaps in the cross-section tables.
@@ 3 Tuning of Pythia 8
[36] | charged particle production | 250 | p, K+p, pp | ||